# DHGS: Decoupled Hybrid Gaussian Splatting for Driving Scene

### v2

@@ Abstract
Existing Gaussian splatting methods often fall short in achieving satisfactory novel view synthesis in driving scenes, primarily due to the absence of crafty design and geometric constraints for the involved elements. This paper introduces a novel neural rendering method termed Decoupled Hybrid Gaussian Splatting (DHGS), targeting at promoting the rendering quality of novel view synthesis for static driving scenes. The novelty of this work lies in the decoupled and hybrid pixel-level blender for road and non-road layers, without the conventional unified differentiable rendering logic for the entire scene, while still maintaining consistent and continuous superimposition through the proposed depth-ordered hybrid rendering strategy. Additionally, an implicit road representation comprised of a Signed Distance Field (SDF) is trained to supervise the road surface with subtle geometric attributes. Accompanied by the use of auxiliary transmittance loss and consistency loss, novel images with imperceptible boundary and elevated fidelity are ultimately obtained. Substantial experiments on the Waymo dataset prove that DHGS outperforms the state-of-the-art methods. The project page where more video evidences are given is: https://ironbrotherstyle.github.io/dhgs_web.
Figure 1: The comparison of rendered road images and ellipsoids on novel view synthesis is presented, with the top row displaying results without SDF regularizer, and the bottom row showcasing results obtained using SDF regularizer. It can be observed that the inclusion of SDF regularization leads the road model to render higher-quality images with the help of enhanced road geometry.
@@ Introduction
Autonomous Driving (AD) is a data-driven task that appeals the attention of 3D reconstruction community for ages. Due to the expense of collecting a tremendous amount of data, the neural rendering technique which mainly consists of Neural Radiance Fields (NeRFs) (Mildenhall et al. 2020; Barron et al. 2021) and Gaussian Splatting (Kerbl et al. 2023) emerges as an alternative. With the gathered AD data, a synthetic driving scenario can be generated, facilitating close-loop simulation and testing, with a variety of sensor configurations not typically achievable in real-world tests Yang et al. (2023b); Tonderski et al. (2024); Lindström et al. (2024). Compared with NeRF-based approaches, researchers utilizing Gaussian splatting can train and deploy a model more efficiently and flexibly. Despite the significant advancements in scene reconstruction offered by Gaussian splatting, there are still imperfections in the novel view synthesis capabilities, especially for critical downstream perceptual tasks such as online mapping (Li et al. 2024) of AD scenes. Current methods either model the whole driving scene in a consolidated manner (Kerbl et al. 2023; Huang et al. 2024a; Cheng et al. 2024; Lu et al. 2024), necessitating a uniform growth and pruning strategy for all scene elements, i.e. road, building, and distant views, or model the near scene and far sky (Wu et al. 2023; Miao et al. 2024) dividually. These approaches tend to emphasize the overall or specific distant elements, often overlooking the nearby synthesis quality, which is brittle and significantly impacted by the change of camera perspective. We believe that prioritizing the road is essential, as its geometric properties are fundamental to the success of AD systems.
@@
In this paper, we propose a static neural rendering method named Decoupled Hybrid Gaussian Splatting for Driving Scene (DHGS) to lift the rendering quality of novel view synthesis, particularly for nearby surroundings with fewer flaws and more subtle details. The insight of the method comes from the existing geometric prior that could be utilized to supervise the nearby road elements along with other non-road environments in a fancy pattern. Specifically, we decouple the whole driving scene into a nearby road model and environment model represented by two distinct Gaussian models, considering optimizing the two models based on their geometric characters accordingly. Before that, point clouds are projected and segmented using semantic 2D masks to derive the initial road and environmental point clouds. The road point cloud is used for pre-training the implicit Signed Distance Field (SDF), which serves as the supervision of road training. The training process is conducted via the introduced depth-ordered hybrid rendering strategy by the blender of rendered images of both models, assisted with transmittance loss that regularizes the accumulated transmittance of each componential Gaussian property, as well as consistency loss which bonds the two areas compactly. Besides, we refine the road elements using an SDF loss that is expressly designed for surface regularization, achieving uniform and orderly road Gaussians. To summarize, the contributions of the proposed DHGS are listed as follows:
@@
    We first propose to decouple the driving scene into near road model and other non-road model, enabling the separate optimization of two models and rendering jointly with the designed depth-ordered hybrid rendering.
@@
    Extensive quantitative and qualitative experiments on the popular Waymo dataset demonstrate that DHGS achieves state-of-the-art performance on scene reconstruction and diversified novel free-view synthesis.
Related Works
@@
NeRF (Mildenhall et al. 2020) has garnered remarkable attention within the neural rendering domain, pioneering the use of a Multi-Layer Perceptron (MLP) to model 3D scenes implicitly. The computationally intensive neural volume rendering and the finite number of sampling points hinder the deployment of NeRF in intricate and expansive driving scenarios. Mip-NeRF (Barron et al. 2021) proposes a more efficient sampling strategy, refining the sampling region from a single ray to a frustum of a viewing cone. Both Mip-NeRF 360 (Barron et al. 2022) and NeRF++ (Zhang et al. 2020) segment the scene into proximal and distal zones, employing warping operations on the latter to handle the unbounded 3D scene. Instant-NGP (Müller et al. 2022) leverages a hash grid for feature storage and streamlines the MLP size to release the computational burden. Plenoxels (Fridovich-Keil et al. 2022) employs sparse grids to reduce storage space and utilize Spherical Harmonics (SH) to represent the appearance. These approaches rely on data from 360-degree object-centric trajectories, while free and long driving scenes draw less attention. F2-NeRF (Wang et al. 2023) proposes a novel space-warping method that supports arbitrary input camera trajectories, enabling the reconstruction of driving scenes. StreetSurf (Guo et al. 2023) extends prior object-centric neural surface reconstruction techniques to address the unique challenges inherent in unbounded street perspectives, demonstrating promise for a range of subsequent downstream tasks. While advancements in ray marching and enhanced MLPs have been notable, these methods remain a trade-off between rendering fidelity and efficiency, which may restrict their utility in expansive autonomous driving scenarios.
@@
3DGS (Kerbl et al. 2023) is a seminal work that pioneers the use of 3D Gaussian ellipsoids to explicitly represent a scene and leverages CUDA for parallel rendering, thereby surpassing NeRFs in both rendering quality and efficiency. HUGS, as introduced in (Zhou et al. 2024a), enables real-time rendering of new perspectives, delivering both 2D and 3D semantic data, delivering both 2D and 3D semantic information with remarkable precision.
@@
Figure 2: The pipeline of the proposed method for driving scene reconstruction. Given consecutive multi-camera images along with their respective road and non-road masks, we initially generate decoupled road pcd (point cloud) and environment pcd, a road SDF is then pre-trained as subsequent guidance for the road Gaussian model. The Environment pcd enables the initialization for the environment Gaussian model which composes rendered images with images by paratactic road model via the proposed depth-ordered hybrid rendering.
@@
The vanilla 3DGS focuses solely on pixel color consistency, often overlooking the modeling of geometric structures. Consequently, obvious discontinuities such as artifacts and holes appear under novel viewpoints. This discontinuity phenomenon would be more prominent on AD datasets with sparse training views. DRGS (Chung, Oh, and Lee 2024) introduces a dense depth map as geometry guidance to mitigate overfitting. NeuSG (Chen, Li, and Lee 2023) flattens the Gaussian ellipsoid by designing a regularization loss for the scale parameter, and performs consistent joint optimization in combination with the signed distance field. Scaffold GS (Lu et al. 2024) precisely controls the density of the Gaussian ellipsoid by establishing voxel anchors and corresponding offset vectors and learning the various parameters of the Gaussian through a simple multi-layer network. Gaussian Pro (Cheng et al. 2024) considers the planar prior in the scene, explicitly constraining the growth of Gaussians, achieving amendatory rendering and more compact representation. 2DGS (Huang et al. 2024a) presents a highly efficient differentiable 2D Gaussian renderer, enabling perspective-accurate splatting by leveraging 2D surface modeling, demonstrating superiority in geometric reconstruction. Driving Gaussians (Zhou et al. 2024b), Street Gaussians (Yan et al. 2024) and S3 Gaussians (Huang et al. 2024b) all initialize the point cloud collected by LiDAR instead of the point cloud generated by SfM. Leveraging the dense, accurate data from LiDAR, additional supervision, such as the positional constraint, is introduced to the learning process.
@@
As far as we know, Gaussian Splatting works for AD in the literature pay little attention to the capability of novel view synthesis, resulting in low-quality images under significant camera transformations. We aim to address this issue by decoupling the entire scene into road and non-road regions, optimizing each with guidance from hybrid rendering and geometric regularization of Gaussians, achieving more robust and higher-fidelity rendering quality.
@@
3DGS utilizes a set of 3D Gaussians to explicitly represent a scene, achieving high-fidelity rendering quality and fast speed via tile-based rasterization. Each Gaussian $G$ is defined as:
$G(\mathbf{x})=e^{-\frac{1}{2}(x-\mu)^{\top}\mathbf{\Sigma}^{-1}(x-\mu)},$ (1)
where $\mu$ is the mean position of Gaussian and $x$ is a discretional coordinate within the Gaussian space. $\mathbf{\Sigma}=RSS^{T}R^{T}$, where $R$ is the rotation matrix and $S$ is the scaling matrix. Apart from the spherical harmonic coefficients that represent the anisotropic color of each Gaussian, each pixel color is aggregated by $\alpha$-blending:
@@
2DGS ensures multi-view consistency well by adopting the expression of surfels and using the projection method specifically for surfels. It achieves 2D Gaussian rendering through ray-splat intersection. The process of ray-splat intersection can be written as:
@@
where $\mathbf{h}_{u}$, $\mathbf{h}_{v}$ are two planes represented in homogeneous coordinates. By defining the function $G(\mathbf{u})=\exp\left(-\frac{u^{2}+v^{2}}{2}\right)$, and employing the object-space low-pass filter $\hat{G}(\mathbf{x})=\max\left\{G(\mathbf{u}(\mathbf{x})),G(\frac{\mathbf{x}-\mathbf{c}}{\sigma})\right\}$, the rasterization for 2DGS can be shown as:
$\mathbf{c}(\mathbf{x})=\sum_{i=1}\mathbf{c}_{i}\alpha_{i}\hat{G}_{i}(\mathbf{u}(\mathbf{x}))\prod_{j=1}^{i-1}(1-\alpha_{j}\hat{G}_{j}(\mathbf{u}(\mathbf{x}))).$ (4)
@@
The proposed method utilizes initial points clouds and semantic masks as paratactic input of multi-camera views. The pcd (point cloud) initialization phase first generates road and non-road point clouds, which will be further modeled as road and environment Gaussian model. Based on the known road point cloud, we propose to design a mathematical implicit road representation by SDF which serves as a prior for the surface training. The SDF-based surface constraint consists of pre-training and off-line supervision phases, by taking advantage of the distance and normal characters geometrically. We use distinct Gaussians to model road and non-road elements, enhancing the rendering quality during perspective shifts. To achieve that, depth-ordered hybrid rendering is delicately designed, by which the road surface and non-road areas can be coupled and superimposed consistently and continuously, yielding superior performance compared with SOTA methods that employ individual Gaussian model. The rendered image by fusing two models will be supervised by ground truth via Gaussian loss, and optimized in collaboration with the regularization terms. We will sequentially detail this process in the following sections.
Initialization of the Decoupled PCD
Distinguishing ourselves from prior methods, we opt to utilize LiDAR scan which exhibits better multi-view consistency and prior geometric structure compared to the point cloud from SfM. The initial point cloud is classified into road and non-road parts for the subsequent training. As shown in Figure 2, based on the calibrated intrinsic and extrinsic parameters, multi-camera images and their masks generated by Mask2Former (Cheng et al. 2022) are utilized to form colored and semantically labeled single-frame point clouds. Then we stitch these separate point clouds to build road and environment point clouds.
@@
To optimize the geometric structure of roads and maintain continuity when facing obvious changes in perspective, we design a road constraint based on surface guidance. Utilizing the segregated road point cloud, we pre-train a signed distance field to represent the road surface implicitly. This approach differs from the joint ray sampling optimization in NeRF and the guidance of geometric structure and image consistency in GSDF (Yu et al. 2024). As road point clouds collected by LiDAR have been calibrated precisely in advance, the excellent geometric structure is suitable to be prior for road learning. A straightforward strategy is directly applying distance constraints to Gaussians that belong to the road. Whereas, considering the spatial geometric characteristics of ellipsoids, imposing constraints only on the centers of Gaussian ellipsoids cannot fully restore their correct geometric shapes, potentially restricting the Gaussian distribution to some extent. We therefore adopt a pre-trained model to optimize the Gaussian distribution by adjusting Gaussian parameters nearing the road surface. To this end, we design specific regularization terms for road-related Gaussians. These include distance and normal direction constraints that ensure Gaussians stay close to the road surface and align with the normal direction of the road surface concurrently.
We utilize a predefined network $f_{\theta}(x)$ for predicting the corresponding SDF value given the positional coordinate of the point cloud. Data normalization is imposed on all the training and testing coordinates to ensure stability and enhanced performance. To equalize the number of points on and off the isosurface, we define the road point cloud as $\mathcal{P}_{r}$, and assign a hypothetical SDF value of zero to each point $x_{p}\in\mathcal{P}_{r}$ on the isosurface. For points off the isosurface, we establish a proximity range around each point in $\mathcal{P}_{r}$, perform random sampling within this range to obtain sample points $\mathcal{P}_{s}\supset\mathcal{P}_{r}$, for $\forall x_{s}\in\mathcal{P}_{s}$, we build a map from point $x_{s}$ to the signed euclidean distance $d_{s}$ as:
@@
where $x_{r}^{s}$ is the nearest point of $x_{s}$ in $\mathcal{P}_{r}$. Considering that bumps exist in uneven road surfaces, we generate the normal direction for each point in $\mathcal{P}_{r}$ to assist in the training of the SDF. To be specific, for $\forall x_{r}\in\mathcal{P}_{r}$, set $X_{r}$ as $k$-nearest neighbors of $x_{r}$, by performing SVD:
@@
We supervise the predicted SDF value from the network with our pre-computed ground truth by incorporating a newly proposed normal loss based on the predicted normal direction. The optimization goal of road SDF is written as:
@@
in which the first and last terms refer to SDF value regularization and eikonal regularizer respectively, the middle term represents normal loss, and ||$\cdot$|| ⋅ | refers to the number of points in the point cloud. The optimal parameters $\theta^{*}=\arg\min_{\theta}{L(\theta)}$ will be frozen and used to guide the learning of road Gaussians in the following processes.
@@
Table 1: The quantitative comparison between our method with four SOTA works on PSNR, SSIM, LPIPS, and FID metrics computed over all clips on the Waymo dataset. We adopt the same training data for each method and follow the corresponding experimental settings strictly. The best and second best are marked in bold and underlined for each metric respectively.
@@
Based on the predefined initial semantic non-road point cloud $\mathcal{P}_{e}$ and road point cloud $\mathcal{P}_{r}$, we build the environment model and road model simultaneously. In this paper, Scaffold GS and 2DGS are chosen as the fundamental Gaussian splatting substrates. By executing forward rendering separately, output images $I_{e}$ and $I_{r}$, as well as depth maps and accumulated transmittance maps will be generated in parallel. In our experiments, simply fusing $I_{e}$ and $I_{r}$ linearly causes inferior image quality. The reason is that naive fusion degrades the learning of Gaussian properties without the solicitude for real depth relationship of Gaussians from the environment and road model. The depth $D$ is rendered based on $d_{i}$ that indicates each splat from the camera, following the rasterization pipeline in (Dai et al. 2024; Chung, Oh, and Lee 2024),
$D=\sum_{i=1}^{N}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),T=\prod_{i=1}^{N}(1-\alpha_{i}).$
Accordingly, we adopt a pixel-level hybrid rendering method based on depth sorting, as depicted in Figure 3. Assuming that the tuple $[I_{e},D_{e},T_{e}]$ severally represents image, depth, and accumulated transmittance rendered by the environment model, $[I_{r},D_{r},T_{r}]$ indicates image, depth, and accumulated transmittance rendered by the road model in the same manner. The composite rendering strategy is presented as:
@@
where $\lambda_{r}$ and $\lambda_{e}$ are the weights for fusing $I_{r}$ and $I_{e}$ rendered from the environment model and road model respectively, which are organized as:
@@
$\begin{matrix}\delta_{D}=\begin{cases}1\hskip 5.69054pt&D_{r}>D_{e}\\ 0&D_{r}\leq D_{e}\end{cases}\end{matrix}$ (10)
One limitation of blending the rendered images discontinuously is that the weights also vary discretely, which will result in the distinct appearance of boundary lines on the edge of road and non-road regions. We thus introduce a consecutive strategy to model the dynamic weights based on the rendered depth. Concretely, the sigmoid function $S(x)$ is utilized to realize the ordered rendering as:
@@
With the proposed smooth rendering mechanism, the overall rendering quality can be significantly improved and the synthesis of new perspectives will appear more natural and fluid.
Figure 4: Comparison of different methods on the Waymo dataset, the left column and right column display the quality of scene reconstruction and novel view synthesis respectively. Our method achieves high-quality reconstruction for both the environment region and the road areas, excelling over other comparative methods in these aspects.
@@
where ${M}$ refers to the semantic mask of the image, ||$\cdot$|| ⋅ | denotes the total number of pixels for given matrix. The $\mathcal{L}_{c}$ enforces consistency between the depth rendered by environment Gaussians and road Gaussians:

### v3

@@ Abstract
Existing Gaussian splatting methods often fall short in achieving satisfactory novel view synthesis in driving scenes, primarily due to the absence of crafty designs and geometric constraints for the involved elements. This paper introduces a novel neural rendering method termed Decoupled Hybrid Gaussian Splatting (DHGS), targeting at promoting the rendering quality of novel view synthesis for static driving scenes. The novelty of this work lies in the decoupled and hybrid pixel-level blender for road and non-road layers, without the conventional unified differentiable rendering logic for the entire scene. Still, consistency and continuity in superimposition are preserved through the proposed depth-ordered hybrid rendering strategy. Additionally, an implicit road representation comprised of a Signed Distance Function (SDF) is trained to supervise the road surface with subtle geometric attributes. Accompanied by the use of auxiliary transmittance loss and consistency loss, novel images with imperceptible boundary and elevated fidelity are ultimately obtained. Substantial experiments on the Waymo dataset prove that DHGS outperforms the state-of-the-art methods. The project page where more video evidences are given is: https://ironbrotherstyle.github.io/dhgs_web.
Figure 1: The comparison of the rendered road images and ellipsoids on novel view synthesis in Waymo dataset, with the top row displaying results without SDF regularizer, and the bottom row showcasing results obtained from the road model using SDF regularizer. It can be observed that the inclusion of SDF regularization leads the road model to render higher-quality images with the help of enhanced road geometry under decoupled scene representation.
@@ Introduction
Autonomous Driving (AD) is a data-driven task that has attracted the attention of the 3D reconstruction community for many years. Due to the expense of collecting a tremendous amount of data, the neural rendering technique which mainly consists of Neural Radiance Fields (NeRFs) (Mildenhall et al. 2020; Barron et al. 2021) and Gaussian Splatting (Kerbl et al. 2023) emerges as an alternative. With the gathered AD data, a synthetic driving scenario can be effectively generated, thereby facilitating closed-loop simulation and testing, with a variety of sensor configurations that are not typically achievable in real-world tests (Yang et al. 2023b; Tonderski et al. 2024; Lindström et al. 2024). Compared with NeRF-based approaches, researchers utilizing Gaussian splatting can train and deploy a model more efficiently and flexibly. Despite the significant advancements in scene reconstruction offered by Gaussian splatting, there are still imperfections in the novel view synthesis capabilities, especially for critical downstream perceptual tasks such as online mapping (Li et al. 2024) of AD scenes. Current methods either model the whole driving scene in a consolidated manner (Kerbl et al. 2023; Huang et al. 2024a; Cheng et al. 2024; Lu et al. 2024), necessitating a uniform growth and pruning strategy for all scene elements, i.e. road, building, and distant views, or model the near scene and far sky (Wu et al. 2023; Miao et al. 2024) dividually. These approaches tend to emphasize the overall or specific distant elements, often overlooking the nearby synthesis quality which is brittle and remarkably impacted by the change of camera perspective. We believe that prioritizing the road is essential, as its geometric properties are fundamental to the success of AD emulation systems.
@@
In this paper, we propose a static neural rendering method named Decoupled Hybrid Gaussian Splatting for Driving Scene (DHGS) to enhance the rendering quality of novel view synthesis, particularly for nearby surroundings with fewer artifacts and more subtle details. The insight of the method comes from the existing geometric prior that could be utilized to supervise the nearby road elements along with other non-road environments in an innovative pattern. Specifically, we decouple the whole driving scene into a nearby road and an environment part represented by two distinct Gaussian models, considering optimizing the two models based on their geometric characters accordingly. Before that, point clouds are projected and segmented using semantic 2D masks to derive the initial road and environment point clouds. The road point cloud is exclusively used for pre-training the implicit Signed Distance Function (SDF), which serves as the supervision of road training. The overall training process is conducted via the introduced depth-ordered hybrid rendering strategy by the blender of rendered images of both models, assisted with transmittance loss that regularizes the accumulated transmittance of each componential Gaussian property, as well as consistency loss which bonds the two areas compactly. Besides, we refine the road elements using the SDF constraint that is expressly designed for the surface regularization, achieving uniform and orderly road Gaussians. As depicted in Figure 1, a robust geometric prior enables the reconstructed scene to retain high-quality structural information even under significant perspective transformations. This capability is of paramount importance for the simulation system to generate hyper-realistic and authentic visual input, which promotes the vehicle to make well-informed decisions in complex driving scenarios. To summarize, the contributions of the proposed DHGS are listed as follows:
@@
    We first propose to decouple the modeling of driving scene into near road model and other non-road model, enabling the separate optimization of two models and rendering jointly with the designed depth-ordered hybrid rendering strategy.
@@
    Extensive quantitative and qualitative experiments on the popular Waymo dataset demonstrate that DHGS achieves the state-of-the-art performance on scene reconstruction and diversified novel free-view synthesis.
Related Work
@@
NeRF (Mildenhall et al. 2020) has garnered remarkable attention within the neural rendering domain, pioneering the use of a Multi-Layer Perceptron (MLP) to model 3D scenes implicitly. The computationally intensive neural volume rendering and the finite number of sampling points hinder the deployment of NeRF in intricate and expansive driving scenarios. Mip-NeRF (Barron et al. 2021) proposes a more efficient sampling strategy, refining the sampling region from a single ray to a frustum of a viewing cone. Both Mip-NeRF 360 (Barron et al. 2022) and NeRF++ (Zhang et al. 2020) segment the scene into proximal and distal zones, employing warping operations on the latter to handle the unbounded 3D scene. Instant-NGP (Müller et al. 2022) leverages a hash grid for feature storage and streamlines the MLP size to release the computational burden. Plenoxels (Fridovich-Keil et al. 2022) employs sparse grids to reduce storage space and utilizes Spherical Harmonics (SH) to represent the appearance. These approaches rely on data from 360-degree object-centric trajectories, while free and long driving scenes draw less attention. F2-NeRF (Wang et al. 2023) proposes a novel space-warping method that supports arbitrary input camera trajectories, enabling the reconstruction of driving scenes. StreetSurf (Guo et al. 2023) extends prior object-centric neural surface reconstruction techniques to address the unique challenges inherent in unbounded street perspectives, demonstrating promise for a range of subsequent downstream tasks. Despite the notable advancements in ray marching and the innovative designs of MLPs, these methods remain a trade-off between rendering fidelity and efficiency, which may restrict their utility in expansive autonomous driving scenarios.
@@
3DGS (Kerbl et al. 2023) is a seminal work that innovates the use of 3D Gaussian ellipsoids to explicitly represent a scene and leverages CUDA for parallel rendering, thus surpassing NeRFs in both rendering quality and efficiency. HUGS (Zhou et al. 2024a) enables real-time rendering of new perspectives, delivering both 2D and 3D semantic information with remarkable precision.
@@
Figure 2: The training pipeline of the proposed method for driving scene reconstruction. Given consecutive multi-camera images along with their respective road and non-road masks, we initially generate decoupled road pcd (point cloud) and environment pcd, a road SDF is then pre-trained as subsequent guidance for the road Gaussian model. The environment pcd enables the initialization for the environment Gaussian model producing $I_{e}$, which is composed to the rendered image $I_{c}$ with image $I_{r}$ from paratactic road model via the proposed depth-ordered hybrid rendering.
@@
The vanilla 3DGS focuses solely on pixel color consistency, commonly overlooking the modeling of geometric structures. Consequently, obvious discontinuities such as artifacts and holes appear under novel viewpoints. This discontinuity phenomenon would be more prominent on AD datasets with sparse training views. DRGS (Chung, Oh, and Lee 2024) introduces the dense depth map as geometry guidance to mitigate overfitting. NeuSG (Chen, Li, and Lee 2023) flattens the Gaussian ellipsoid by designing a regularization loss for the scale parameter, and performs consistent joint optimization in combination with the signed distance function. Scaffold GS (Lu et al. 2024) precisely controls the density of the Gaussian ellipsoids by establishing voxel anchors and corresponding offset vectors and learns the various parameters of the Gaussian through a simple multi-layer network. Gaussian Pro (Cheng et al. 2024) considers the planar prior in the scene, explicitly constraining the growth of Gaussians, achieving amendatory rendering and more compact representation. 2DGS (Huang et al. 2024a) presents a highly efficient differentiable 2D Gaussian renderer, enabling perspective-accurate splatting by leveraging 2D surface modeling, demonstrating superiority in geometric reconstruction. Driving Gaussians (Zhou et al. 2024b), Street Gaussians (Yan et al. 2024) and S3 Gaussians (Huang et al. 2024b) all initialize the point cloud collected by the LiDAR instead of the point cloud generated by Structure from Motion (SfM). By leveraging dense and accurate data from LiDAR, additional supervision, such as the positional constraint, is introduced to the learning process.
@@
As far as we know, current Gaussian Splatting works for AD in the literature pay little attention to the capability of novel view synthesis, resulting in low-quality images under significant camera transformations. We aim to address this issue by decoupling the entire scene into road and non-road regions, optimizing each with the guidance from hybrid rendering and geometric regularization of Gaussians, achieving more robust and higher-fidelity rendering quality.
@@
3DGS (Kerbl et al. 2023) utilizes a set of 3D Gaussians to explicitly represent a scene, achieving high-fidelity rendering quality and fast speed via tile-based rasterization. Each Gaussian $G$ is defined as:
$G({x})=e^{-\frac{1}{2}(x-\mu)^{\top}{\Sigma}^{-1}(x-\mu)},$ (1)
where $\mu$ is the mean position of Gaussian and $x$ is a discretional coordinate within the Gaussian space. ${\Sigma}=RSS^{T}R^{T}$, where $R$ is the rotation matrix and $S$ is the scaling matrix. Apart from the spherical harmonic coefficients that represent the anisotropic color of each Gaussian, the color of each pixel $\mathbf{x}$ is aggregated by $\alpha$-blending:
@@
2DGS (Huang et al. 2024a) ensures multi-view consistency well by adopting the expression of surfels and using the projection method specifically for surfels. It achieves 2D Gaussian rendering through ray-splat intersection. The process of ray-splat intersection can be written as:
@@
where $\mathbf{h}_{u}$, $\mathbf{h}_{v}$ are two planes represented in homogeneous coordinates. For the point $\mathbf{u}=(u,v)$, defining the function $G(\mathbf{u})=\exp\left(-\frac{u^{2}+v^{2}}{2}\right)$, and employing the object-space low-pass filter $\hat{G}(\mathbf{x})=\max\left\{G(\mathbf{u}(\mathbf{x})),G(\frac{\mathbf{x}-\mathbf{c}}{\sigma})\right\}$, the rasterization for 2DGS can be shown as:
$\mathbf{c}(\mathbf{x})=\sum_{i=1}{c}_{i}\alpha_{i}\hat{G}_{i}(\mathbf{u}(\mathbf{x}))\prod_{j=1}^{i-1}(1-\alpha_{j}\hat{G}_{j}(\mathbf{u}(\mathbf{x}))).$ (4)
@@
We illustrate the overview of the framework in Figure 2. The proposed method utilizes initial points clouds and semantic masks as paratactic inputs of multi-camera views. The pcd initialization phase first generates road and non-road point clouds, which will be further modeled as decoupled road and environment Gaussian models. Based on the known road point cloud, we propose to design a neural implicit road representation using SDF which serves as a prior for the surface training. The SDF-based surface constraint consists of pre-training and offline supervision phases, by taking advantage of the distance and normal characteristics geometrically. We use distinct Gaussians to model road and non-road elements, enhancing the rendering quality during perspective shifts. To achieve that, the depth-ordered hybrid rendering is delicately designed, by which the road surface and non-road areas can be coupled and superimposed consistently and continuously, yielding superior performance compared with SOTA methods that employ individual Gaussian model. The rendered image resulting from the hybrid fusion of two models is supervised by ground truth via Gaussian loss and optimized along with the regularization terms. We will sequentially detail these processes in the following sections.
Initialization of the Decoupled PCDs
For a given clip, we opt to utilize point cloud from the LiDAR that exhibits better multi-view consistency and prior geometric structure compared to the point cloud from SfM. The initial point cloud is divided into road and non-road parts for the subsequent training. As shown in Figure 2, based on the calibrated intrinsic and extrinsic parameters, multi-camera images and their masks generated by Mask2Former (Cheng et al. 2022) are utilized to form colored and semantically labeled single-frame point clouds. These individual point clouds are then stitched together to construct road and environment point clouds that cover the entire sequence.
@@
To optimize the geometric structure of the road and maintain continuity when facing obvious changes in perspective, we design a road constraint based on the surface guidance. Utilizing the segregated road point cloud, we pre-train a signed distance function to represent the road surface implicitly. This approach differs from the joint ray sampling optimization in NeRF and the guidance of geometric structure and image consistency in GSDF (Yu et al. 2024). As road point clouds collected by LiDAR have been calibrated precisely in advance, the excellent geometric structure is suitable to be the prior for road learning. A straightforward strategy is directly applying distance constraints to Gaussians that belong to the road. Whereas, considering the spatial geometric characteristics of ellipsoids, imposing constraints only on the centers of Gaussian ellipsoids cannot fully restore their correct geometric shapes, potentially restricting the distribution of Gaussian to some extent. We therefore adopt a pre-trained model to optimize the distribution of Gaussian by adjusting Gaussian parameters nearing the road surface. To this end, we design specific regularization terms for road-related Gaussians. These include distance and normal direction constraints that ensure Gaussians stay close to the road surface and align with the normal direction of the road surface concurrently.
We utilize a predefined network $f_{\theta}(x)$ for predicting the corresponding SDF value given the positional coordinate of the point cloud. Data normalization is imposed on all the training and testing coordinates to ensure stability and elevated performance. To equalize the number of points on and off the isosurface, we define the road point cloud as $\mathcal{P}_{r}$, and assign a hypothetical SDF value of zero to each point $x_{p}\in\mathcal{P}_{r}$ on the isosurface. For points off the isosurface, we establish a proximity range around each point in $\mathcal{P}_{r}$, and perform random sampling within this range to obtain sample points $\mathcal{P}_{s}\supset\mathcal{P}_{r}$. For $\forall x_{s}\in\mathcal{P}_{s}$, we build a map from point $x_{s}$ to the signed euclidean distance $d_{s}$ as:
@@
where $x_{r}^{s}$ is the nearest point of $x_{s}$ in $\mathcal{P}_{r}$. Considering that bumps exist in uneven road surfaces, we generate the normal direction for each point in $\mathcal{P}_{r}$ to assist in the training of the SDF. To be specific, for $\forall x_{r}\in\mathcal{P}_{r}$, set $X_{r}$ as $k$-nearest neighbors of $x_{r}$, by performing Singular Value Decomposition (SVD):
@@
We supervise the predicted SDF value from the network with the pre-computed ground truth by incorporating a newly proposed normal loss based on the predicted normal direction. The optimization goal for road SDF is written as:
@@
in which the first and last terms refer to SDF value loss and eikonal loss respectively, the middle term represents the normal loss, and ||$\cdot$|| ⋅ | refers to the number of points in the point cloud. The optimal parameters $\theta^{*}=\arg\min_{\theta}{L(\theta)}$ will be frozen and used to guide the learning of road Gaussians in the following processes.
@@
Table 1: The quantitative comparison between our method with four SOTA works on PSNR, SSIM, LPIPS, and FID metrics computed over all clips on the Waymo dataset. We adopt the same training data for each method and follow the corresponding experimental settings strictly. The details of Set1 to Set4 are described in the appendix. The best and second best are marked in bold and underlined for each metric respectively.
@@
Based on the predefined initial semantic non-road point cloud $\mathcal{P}_{e}$ and road point cloud $\mathcal{P}_{r}$, we build the environment model and road model simultaneously. In this paper, Scaffold GS and 2DGS are chosen as the fundamental Gaussian splatting substrates. By executing forward rendering separately, output images $I_{e}$ and $I_{r}$ are generated separately, along with depth maps and accumulated transmittance maps simultaneously. The depth map $D$ is rendered based on $d_{i}$ that indicates each splat from the camera, following the rasterization pipeline in (Dai et al. 2024; Chung, Oh, and Lee 2024). We give the definitions of $D$ and $T$ as:
$D=\sum_{i=1}^{N}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),T=\prod_{i=1}^{N}(1-\alpha_{i}).$ (8)
In our experiments, linearly fusing $I_{e}$ and $I_{r}$ results in inferior image quality. The reason is that the simplistic fusion overlooks the learning of Gaussian properties and the actual depth relationships of Gaussians derived from the environment and road model. Accordingly, we adopt a pixel-level hybrid rendering method based on depth sorting, as depicted in Figure 3. Assuming that the tuple $[I_{e},D_{e},T_{e}]$ severally represents the image, depth, and accumulated transmittance rendered by the environment model, $[I_{r},D_{r},T_{r}]$ indicates the image, depth, and accumulated transmittance rendered by the road model in the same manner. The composite rendering strategy is presented as:
@@
where $\lambda_{r}$ and $\lambda_{e}$ are the weights for fusing $I_{r}$ and $I_{e}$, which are organized as:
@@
$\begin{matrix}\delta_{D}=\begin{cases}1\hskip 5.69054pt&D_{r}>D_{e},\\ 0&D_{r}\leq D_{e}.\end{cases}\end{matrix}$ (11)
One limitation of blending the rendered images discontinuously is that the weights also vary discretely, which results in the distinct appearance of boundary lines on the edge of road and non-road regions. We thus introduce a consecutive strategy to model the dynamic weights based on the rendered depth. Concretely, the sigmoid function $S(x)$ is utilized to realize the ordered rendering as:
@@
With the proposed smooth rendering mechanism, the overall rendering quality can be considerably improved, and the synthesis of new perspectives will appear more natural and fluid.
Figure 4: Comparison of different methods on the Waymo dataset, the left column and right column display the quality of scene reconstruction and novel view synthesis respectively. Our method achieves high-quality reconstruction for both the environment and the road areas, excelling over other comparative methods.
@@
where ${M}$ refers to the semantic mask of the image, with entries of 1 for road pixels and 0 for non-road pixels. ||$\cdot$|| ⋅ | denotes the total number of pixels for a given matrix. The $\mathcal{L}_{\text{cons}}$ enforces consistency between the depth rendered by environment Gaussians and road Gaussians: